OpenStudy (anonymous):

What is the volume of a cylinder with a radius of 7 feet and a height of 12 feet? Use 3.14 for pi. Round your answer to the nearest hundredth. 1,846.32 cubic feet 1,950.25 cubic feet 1,999.21 cubic feet 2,000.31 cubic feet

OpenStudy (anonymous):

@P0sitr0n

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

volume = (area of base) * (height) area of base = pi (7)^2 height = 12 volume = 49 (3.14) (12) = ?